A man makes 2.5 laps around a circular lap 2.5 minutes. The circumference of the track is 126 m while its diameter is 40 m. Compute the man’s (a) average speed, and (b) average velocity.


1
Expert's answer
2021-10-19T15:10:31-0400

(a) By the definition of the average speed we have:


"v_{avg}=\dfrac{Total\ distance}{Total\ elapsed\ time}=\dfrac{2.5\times C}{t},""v_{avg}=\dfrac{2.5\times126\ m}{2.5\ min\times\dfrac{60\ s}{1\ min}}=2.1\ \dfrac{m}{s}."

(b) By the definition of the average velocity we have:


"\vec{v}_{avg}=\dfrac{Displacement}{Total\ time}."

In our case displacement is 0.5 lap which is equal to the diameter of the track. Therefore, we get:


"\vec{v}_{avg}=\dfrac{40\ m}{2.5\ min\times\dfrac{60\ s}{1\ min}}=0.27\ \dfrac{m}{s}."